# To See is Not to Master: Teaching LLMs to Use Private Libraries for Code Generation

### 19 Mar 2026 v3

@@ 1. Introduction
One fundamental reason why LLMs perform poorly on private libraries is that such libraries are typically absent from the training corpora (Zan et al., 2022). This naturally suggests training LLMs on data related to the target private library. A natural next question, then, is how to obtain training data about private libraries at scale. Due to the closed-source nature of private libraries, real-world code that invokes them is often scarce, while manually curating such data is labor-intensive and impractical (Li et al., 2024b; Majumdar et al., 2025). We therefore turn to LLM-driven data synthesis, where LLMs automatically generate training samples that are subsequently used to fine-tune themself (Luo et al., 2023; Liu et al., 2026; Wei et al., 2023; Wu et al., 2025). However, direct synthesis is far from trivial, because the models initially lack sufficient prior knowledge of the target private library. As a result, directly synthesized data often suffers from two major challenges. ❶ Low Data Diversity. LLMs tend to generate overly basic requirements that can be solved with only a small number of private API. In contrast, real-world development scenarios often require the coordinated invocation of multiple private APIs to satisfy diverse and complex coding requirements. This creates a substantial gap between synthesized data and practical development needs. ❷ Poor Data Quality. LLMs are prone to producing flawed training samples, such as those containing syntax errors, non-existent APIs, or semantically incorrect API invocations. Directly training on such noisy data can be counterproductive and may even degrade the model’s overall capabilities.
@@
Evaluating private-library-oriented code generation requires libraries that are unseen to the model. However, many existing benchmarks are built on libraries that have been publicly released for a long time, making them unreliable for assessing the private-library-oriented generation capability of today’s popular LLMs (Zan et al., 2022; Wang et al., 2025b). To obtain a more faithful evaluation, we select two recently released libraries, ndonnx (QuantCo, 2025) and numba-cuda (NVIDIA, 2026), both of which were introduced in 2024 and continued to evolve throughout 2025. Based on them, we carefully construct two new benchmarks, namely NdonnxEval and NumbaEval, which contain 169 and 187 instances, respectively. We apply PriCoder to three mainstream LLMs. Our results show that PriCoder substantially improves LLMs’ private-library-oriented code generation capability, yielding gains of more than 20% in pass@1 in many settings. Additionally, these gains come with negligible impact on the models’ general capabilities. Ablation studies further confirm that both Progressive Graph Evolution and Multidimensional Graph Pruning are essential to the overall effectiveness of our approach.
@@ 3. Motivation
Most prior work on private-library-oriented code generation is based on RAG. They implicitly assume that once the model is presented with the required API knowledge, it can reliably invoke required APIs to satisfy the coding requirement. In this section, we challenge this assumption and show that seeing the right API knowledge does not necessarily translate into invoking it correctly.
@@
We further analyze the failed cases and find that errors are largely attributable to ineffective invocation of the provided oracle APIs, rather than missing any required knowledge in the context. In many instances, the model either does not invoke a necessary API even when it is explicitly available, or invokes it in an incorrect manner. We highlight two representative cases in Figure 1. ➀ In the first case, the task requires converting an array into an ndonnx.Array via ndonnx.asarray before applying ndonnx.multiply. Despite being provided with both API specifications, the model directly calls ndonnx.multiply on the raw input and omits the required conversion, which triggers a type error at runtime. ➁ In the second case, the task requires safely handling division-by-zero by using ndonnx.where to select between the computed quotient and a fallback value. Although the model attempts to use ndonnx.where, it misinterprets the API signature and calls ndonnx.where with only one argument, leading to an argument error.
@@
Motivated by the above findings, we argue that LLMs should be trained to learn how to invoke private-library APIs. However, training data for private libraries is highly scarce. We therefore consider a natural direction: synthesizing training data with LLMs (Majumdar et al., 2025; Adarsh et al., 2025). However, naively prompting LLMs to directly generate such training data about private libraries faces some critical challenges. To illustrate this, we conduct an additional study on ndonnx, a library that is unfamiliar to all evaluated models in this paper. Specifically, we inject the complete API documentation of the library into the prompt and ask DeepSeek-Coder-6.7B-Instruct to directly synthesize $200$ training instances, each consisting of a coding requirement and its corresponding solution.
@@ 4.1. Overview
Starting from API nodes, PriCoder builds $\mathcal{G}$ into a larger graph step by step. As discussed in Section 3, naively synthesized data often suffers from low diversity and poor quality. To address this, PriCoder introduces two graph operators. First, Progressive Graph Evolution (Section 4.2) expands the graph by synthesizing new sample nodes based on existing nodes, enabling the graph to grow from API-level knowledge to increasingly diverse training samples. Second, Multidimensional Graph Pruning (Section 4.3) removes low-quality sample nodes through automatic verification, ensuring that only reliable nodes remain in the graph. By alternately evolving and pruning the graph, PriCoder automatically constructs a high-diversity and high-quality synthetic dataset $\mathcal{D}_{\text{syn}}=\{(r_{i},y_{i})\}_{i=1}^{N}$ from the retained sample nodes. Figure 2 provides an overview of PriCoder.
@@ 4.2. Progressive Graph Evolution
To address the low-diversity challenge identified in Section 3, we introduce Progressive Graph Evolution (Algorithm 2), a graph operator that progressively grows the synthesis graph from basic nodes to increasingly diverse ones. Specifically, we first evolve the graph from API nodes into initial sample nodes, and then iteratively expand it by synthesizing new sample nodes based on existing sample nodes. In this way, the graph can explore a much larger API composition space, thereby producing increasingly diverse training samples.
@@ 4.3. Multidimensional Graph Pruning
To address the poor-quality challenge identified in Section 3, we introduce Multidimensional Graph Pruning (Algorithm 3), a graph operator that removes low-quality sample nodes through multidimensional quality verification. For any new sample node $(r,y)$, we retain it in the Graph only if it passes three complementary checks which target different failure modes: syntactic correctness, execution validity, and overall functionality. Formally, we define three binary validators $V_{\text{syn}},V_{\text{exe}},V_{\text{sem}}\in\{0,1\}$ and retain a sample node in the Graph only if
(4) $V(r,y)=V_{\text{syn}}(r,y)\cdot V_{\text{exe}}(r,y)\cdot V_{\text{sem}}(r,y)=1.$
@@ 5. Benchmark Construction
Evaluating private-library-oriented code generation requires benchmarks built on target libraries that are unseen by the tested models. However, many existing benchmarks rely on libraries that were released long ago and may have already been fully learned by recent LLMs. For example, TorchDataEval (Zan et al., 2022), a benchmark widely used in prior work (Zan et al., 2022, 2024, 2025), evaluates models on the TorchData library (Meta, 2021), which was released in December 2021, well before the knowledge cutoffs of most modern LLMs. As a result, improvements on such benchmarks do not necessarily provide convincing evidence that a method truly enhances private-library-oriented code generation.
@@
LLM-Driven Draft Generation. To reduce manual effort, we leverage GPT-5.2 (Singh et al., 2025) to generate initial benchmark drafts, as it has already acquired sufficient knowledge of both libraries. For each library, we prompt the model to generate more than 500 draft instances, each consisting of a coding requirement and its corresponding unit tests.
Manual Curation and Refinement. To ensure benchmark quality, five student authors with extensive programming experience (more than four years on average) manually filtered and refined the generated drafts. To further ensure that each instance is indeed solvable, we also provided a reference solution for every instance and verified that it passes all corresponding test cases. After more than 150 person-hours of collective effort, we finalized 169 instances for NdonnxEval and 187 instances for NumbaEval. On average, each instance contains more than 9 test cases and requires the coordinated use of more than 4 distinct APIs. Detailed statistics of the two benchmarks are summarized in Table 2.
@@ 6.1. Research Questions
RQ4: How key factors affects the effectiveness of PriCoder? This RQ studies how key factors influence the effectiveness of PriCoder. In particular, we focus on two key factors: the scale of the synthesized training data and the LLM used for data synthesis, and analyze how they affect the performance of PriCoder.
@@ 6.2. Benchmarks
  • NumbaEval and NdonnxEval (Section 5). We release two novel benchmarks for rigorous evaluation of private-library-oriented code generation. NumbaEval evaluates code generation with numba-cuda library, and NdonnxEval evaluates code generation with ndonnx library.
@@ 6.3. Metrics
We adopt unbiased pass@k and exec@k as our main metrics.
@@ 6.4. Baselines
  • Naive RAG (Zhou et al., 2022; Zan et al., 2022). It adopts a straightforward retrieval-augmented generation pipeline to search for relevant APIs and include their specifications directly in the prompt.
@@ 6.6. Implementation Details
For baselines that require task decomposition or reranking, we use the same LLM as the backbone for these steps. We adopt the widely used BGE-M3 (Multi-Granularity, 2024) as the embedding model for retrieval. For baseline-specific hyperparameters, we perform grid search on NdonnxEval and use the best-performing configuration.
To ensure fair comparison, all methods, including PriCoder, generate 10 samples per requirement with identical decoding settings, using temperature 0.5 and top-$p$ 0.95. We use vLLM (Kwon et al., 2023) as the inference framework and LLaMA-Factory (Zheng et al., 2024) for model fine-tuning. All experiments are conducted on a server equipped with eight NVIDIA A100 GPUs, each with 80 GB of memory. Additional details are provided in the Supplementary Material.
@@ 7.2. RQ2: Impact on General Code Generation
PriCoder can even enhance a model’s general code generation capabilities. Interestingly, in several settings, PriCoder actually improves both pass@k and exec@k. For instance, PriCoder increases the pass@5 of LLaMa from 78.17% to 78.94% and its exec@1 from 95.37% to 96.46%. We attribute this unexpected benefit to the rigorous Multidimensional Graph Pruning. By exclusively training on data that has passed Multidimensional Graph Pruning, the model reinforces its adherence to correct syntax and programming paradigms. As shown in Figure 3, the Vanilla model fails to use isinstance, whereas the model trained with PriCoder applies it correctly and successfully passes the test cases.
@@ 7.3. RQ3: Ablation Study
PriCoder consists of two key components, namely Progressive Graph Evolution and Multidimensional Graph Pruning. In this RQ, we conduct an ablation study to quantify the contribution of each component to the overall effectiveness of PriCoder.
@@
Results. Table 5 reports the ablation results on NdonnxEval and NumbaEval, and Table 6 reports the results on HumanEval.
Both Progressive Graph Evolution and Multidimensional Graph Pruning contribute to private-library-oriented code generation. Removing either component leads to a clear performance drop on both NdonnxEval and NumbaEval. For example, without Progressive Graph Evolution, pass@5 on NdonnxEval decreases from 62.70% to 57.66%. Without Multidimensional Graph Pruning, pass@5 on NumbaEval drops from 58.29% to 46.92%. The performance degradation becomes even more severe under direct synthesis, where exec@1 on NumbaEval further drops sharply from 56.47% to 31.55%.
Multidimensional Graph Pruning helps preserve general code generation capability. When Multidimensional Graph Pruning is removed, both pass@k and exec@k on HumanEval decline. This proved that training on unfiltered synthesized data can introduce harmful noise, which may reinforce incorrect coding patterns and negatively affect general code generation.
@@ 7.4. RQ4: Analyzing Data Synthesis in PriCoder
Impact of the scale of the synthesized data.Performance improves steadily as the scale of synthesized data increases. As the training set grows, the model achieves progressively better results. For example, when the data scale increases from 10,000 to 40,000, pass@5 improves from 61.5% to 68.5%. This trend suggests that organizations with sufficient resources can further strengthen private-library-oriented code generation by scaling up synthesized training data. ❷ Substantial gains can be achieved even with a small amount of synthesized data. Even with only 1,000 synthesized training samples, pass@5 and exec@5 increase from 33.3% and 42.2% to 50.6% and 61.9%, respectively. This indicates that PriCoder can already deliver meaningful improvements with a relatively small training set, making it practical for frequent and low-cost adaptation to newly evolved private-library knowledge.

### 27 Mar 2026 v4

@@ 1. Introduction
One fundamental reason why LLMs perform poorly on private libraries is that such libraries are typically absent from the training corpora (Zan et al., 2022). This naturally suggests training LLMs on data related to the target private library. A natural next question, then, is how to obtain training data about private libraries at scale. Due to the closed-source nature of private libraries, real-world code that invokes them is often scarce, while manually curating such data is labor-intensive and impractical (Li et al., 2024b; Majumdar et al., 2025). We therefore turn to LLM-driven data synthesis, where LLMs automatically generate training samples that are subsequently used to fine-tune themselves (Luo et al., 2023; Liu et al., 2026; Wei et al., 2023; Wu et al., 2025). However, direct synthesis is far from trivial. As shown in Section 7.3, fine-tuning models on directly synthesized data yields a marginal improvement of less than 10% in pass@1. This limited performance gain is primarily because models initially lack sufficient prior knowledge of the target private library, causing the directly synthesized data to suffer from two major challenges. ❶ Low Data Diversity. LLMs tend to generate overly basic requirements that can be solved with only a small number of private APIs. In contrast, real-world development scenarios often require the coordinated invocation of multiple private APIs to satisfy diverse and complex coding requirements. This creates a substantial gap between synthesized data and practical development needs. ❷ Poor Data Quality. LLMs are prone to producing flawed training samples, such as those containing syntax errors, non-existent APIs, or semantically incorrect API invocations. Directly training on such noisy data can be counterproductive and may even degrade the model’s overall capabilities.
@@
Evaluating private-library-oriented code generation requires libraries that are unseen to the model. However, many existing benchmarks are built on libraries that have been publicly released for a long time, making them unreliable for assessing the private-library-oriented generation capability of today’s popular LLMs (Zan et al., 2022; Wang et al., 2025b). To obtain a more faithful evaluation, we select two recently released libraries, ndonnx (QuantCo, 2025) and numba-cuda (NVIDIA, 2026), both of which were introduced in 2024 and continued to evolve throughout 2025. Based on them, we carefully construct two new benchmarks, namely NdonnxEval and NumbaEval, which contain 169 and 187 instances, respectively. We apply PriCoder to three mainstream LLMs. Our results show that PriCoder substantially improves LLMs’ private-library-oriented code generation capability, yielding gains of more than 30% in pass@k in many settings. Additionally, these gains come with negligible impact on the models’ general capabilities. Ablation studies further confirm that both Progressive Graph Evolution and Multidimensional Graph Pruning are essential to the overall effectiveness of our approach.
@@ 3. Motivation
Most prior work on private-library-oriented code generation is based on RAG. They implicitly assume that once the model is presented with the required knowledge, it can reliably invoke required APIs to satisfy the coding requirement. In this section, we challenge this assumption and show that seeing the sufficient required API knowledge does not necessarily translate into invoking it correctly.
@@
We further analyze the failed cases and find that errors are largely attributable to ineffective invocation of the provided oracle APIs, rather than missing any required knowledge in the context. In many instances, the model either does not invoke a necessary API even when it is explicitly required, or invokes it in an incorrect manner. We highlight two representative cases in Figure 1. ➀ In the first case, the task requires converting an array into an ndonnx.Array via ndonnx.asarray before applying ndonnx.multiply. Despite being provided with both API specifications, the model directly calls ndonnx.multiply on the raw input and omits the required conversion, which triggers a type error at runtime. ➁ In the second case, the task requires safely handling division-by-zero by using ndonnx.where to select between the computed quotient and a fallback value. Although the model attempts to use ndonnx.where, it misinterprets the API signature and calls ndonnx.where with only one argument, leading to an argument error.
@@
Motivated by the above findings, we argue that LLMs should be trained to learn how to invoke private-library APIs. However, training data for private libraries is highly scarce. We therefore consider a natural direction: synthesizing training data with LLMs (Majumdar et al., 2025; Adarsh et al., 2025). However, naively prompting LLMs to directly generate such training data about private libraries faces some critical challenges. To illustrate this, we conduct an additional empirical study on ndonnx, a library that is unfamiliar to all evaluated models in this paper. Specifically, we inject the complete API documentation of the library into the prompt and ask DeepSeek-Coder-6.7B-Instruct to directly synthesize $200$ training instances, each consisting of a coding requirement and its corresponding solution.
@@ 4.1. Overview
Starting from API nodes, PriCoder builds $\mathcal{G}$ into a larger graph step by step. As discussed in Section 3, naively synthesized data often suffers from low diversity and poor quality. To address these challenges, PriCoder introduces two graph operators. First, Progressive Graph Evolution (Section 4.2) expands the graph by synthesizing new sample nodes based on existing nodes, enabling the graph to grow from API-level knowledge to increasingly diverse training samples. Second, Multidimensional Graph Pruning (Section 4.3) removes low-quality sample nodes through automatic verification, ensuring that only reliable nodes remain in the graph. By alternately evolving and pruning the graph, PriCoder automatically constructs a high-diversity and high-quality synthetic dataset $\mathcal{D}_{\text{syn}}=\{(r_{i},y_{i})\}_{i=1}^{N}$ from the retained sample nodes. Figure 2 provides an overview of them.
@@ 4.2. Progressive Graph Evolution
To address the low-diversity challenge identified in Section 3, we introduce Progressive Graph Evolution (Algorithm 2), a graph operator that progressively grows the synthesis graph from basic nodes to increasingly diverse ones. Specifically, we first evolve the graph from API nodes into initial sample nodes, and then expand it by synthesizing new iterative sample nodes based on existing sample nodes. In this way, the graph can explore a much larger API composition space, thereby producing increasingly diverse samples.
@@ 4.3. Multidimensional Graph Pruning
To address the poor-quality challenge identified in Section 3, we introduce Multidimensional Graph Pruning (Algorithm 3), a graph operator that removes low-quality sample nodes through multidimensional quality verification. For any new sample node $(r,y)$, we retain it in the graph only if it passes three complementary checks which target different failure modes: syntactic correctness, execution validity, and overall functionality. Formally, we define three binary validators $V_{\text{syn}},V_{\text{exe}},V_{\text{fun}}\in\{0,1\}$ and retain a sample node in the graph only if
(4) $V(r,y)=V_{\text{syn}}(r,y)\cdot V_{\text{exe}}(r,y)\cdot V_{\text{fun}}(r,y)=1.$
@@ 5. Benchmark Construction
Evaluating private-library-oriented code generation requires target libraries unseen by the tested models. However, many existing benchmarks are built on libraries released long ago and may already be familiar to recent LLMs. For example, TorchDataEval (Zan et al., 2022), a benchmark widely used in prior work (Zan et al., 2022, 2024, 2025), evaluates models on the TorchData library (Meta, 2021), which was released in December 2021, well before the knowledge cutoffs of most modern LLMs. Therefore, gains on such benchmarks do not necessarily provide convincing evidence of improved private-library-oriented code generation.
@@
Instance Construction. For each library, five annotators with extensive programming experience (more than four years on average) manually constructed benchmark instances, where each instance consists of a coding requirement and its corresponding unit tests. To improve construction efficiency, the annotators were allowed to use GPT-5.2 (Singh et al., 2025) as an auxiliary tool for documentation consultation during the construction process.
Verification and Refinement. To ensure benchmark quality, all constructed instances were further manually verified and refined by the annotators. In particular, the annotators carefully checked the correctness, clarity, and realism of each requirement and its associated test cases. To further ensure that each instance is indeed solvable, we also provided a reference solution for every instance and verified that it passes all corresponding test cases.
After more than 150 person-hours of collective effort, we finalized 169 instances for NdonnxEval and 187 instances for NumbaEval. On average, each instance contains more than 9 test cases and requires the coordinated use of more than 4 distinct APIs. Detailed statistics of the two benchmarks are summarized in Table 2.
@@ 6.1. Research Questions
RQ4: How do key factors affect the effectiveness of PriCoder? This RQ studies how key factors influence the effectiveness of PriCoder. In particular, we focus on two key factors: the scale of the synthesized training data and the LLM used for data synthesis, and analyze how they affect the performance of PriCoder.
@@ 6.2. Benchmarks
NumbaEval and NdonnxEval (Section 5). We introduce two novel benchmarks for rigorous evaluation of private-library-oriented code generation. NumbaEval is built on the numba-cuda library, while NdonnxEval is built on the ndonnx library.
@@ 6.3. Metrics
We use pass@k and exec@k as our main evaluation metrics. To reduce the randomness and obtain more reliable estimates, we compute both metrics using the standard unbiased estimator.
@@ 6.4. Baselines
  • Naive RAG (Zhou et al., 2022; Zan et al., 2022). It adopts a retrieval-augmented generation pipeline to search for relevant APIs and include their specifications directly in the prompt.
@@ 6.6. Implementation Details
For baselines that require task decomposition or reranking, we use the same LLM as the backbone for these steps. We adopt the widely used BGE-M3 (Multi-Granularity, 2024) as the embedding model for retrieval. To ensure fair comparison, all methods, including PriCoder, generate 10 samples per requirement with identical decoding settings, using temperature 0.5 and top-$p$ 0.95. We use vLLM (Kwon et al., 2023) as the inference framework and LLaMA-Factory (Zheng et al., 2024) for model fine-tuning. All experiments are conducted on a server equipped with eight NVIDIA A100 GPUs, each with 80 GB of memory. Additional details are provided in the Supplementary Material.
@@ 7.2. RQ2: Impact on General Code Generation
PriCoder can even enhance a model’s general code generation capabilities. Interestingly, in several settings, PriCoder actually improves both pass@k and exec@k. For instance, PriCoder increases the pass@5 of LLaMa-8B from 78.17% to 78.94% and its exec@1 from 95.37% to 96.46%. We attribute this unexpected benefit to the Multidimensional Graph Pruning. By exclusively training on data that has passed Multidimensional Graph Pruning, the model reinforces its adherence to correct syntax and programming paradigms. As shown in Figure 3, the Vanilla model fails to use isinstance, whereas the model trained with PriCoder applies it correctly and successfully passes the test cases.
@@ 7.3. RQ3: Ablation Study
PriCoder consists of two key components. In this RQ, we conduct an ablation study to quantify the contribution of each component to the overall effectiveness of PriCoder.
@@
Results. The results are shown in Table 5 and Table 6.
Both Progressive Graph Evolution and Multidimensional Graph Pruning contribute to private-library-oriented code generation. Removing either component leads to a clear performance drop on both NdonnxEval and NumbaEval. For example, without Progressive Graph Evolution, pass@5 on NdonnxEval decreases from 62.76% to 57.66%. Without Multidimensional Graph Pruning, pass@5 on NumbaEval drops from 58.29% to 46.92%. The performance degradation becomes even more severe under direct synthesis, where exec@1 on NumbaEval further drops sharply from 56.47% to 31.55%.
Multidimensional Graph Pruning helps preserve general code generation capability. When Multidimensional Graph Pruning is removed, both pass@k and exec@k on HumanEval decline. This proves that training on unfiltered synthesized data can introduce harmful noise, which may reinforce incorrect coding patterns and negatively affect general code generation.
@@ 7.4. RQ4: Analyzing Data Synthesis in PriCoder
Impact of the scale of synthesized data.Performance improves steadily as synthesized data scales up. As the training set grows, the model achieves progressively better results. For example, increasing the data scale from 10,000 to 40,000 raises pass@5 from 61.5% to 68.5%. This suggests that, with sufficient resources, organizations can further improve private-library-oriented code generation by scaling synthesized training data. ❷ Substantial gains can be achieved even with a small amount of synthesized data. With only 1,000 synthesized training samples, pass@5 and exec@5 increase from 33.3% and 42.2% to 50.6% and 61.9%, respectively. This shows that PriCoder can already provide substantial improvements with a relatively small training set, making it practical for low-cost adaptation to newly evolved private-library knowledge.
@@
8. Discussion
8.1. Reliability of LLM Judgment
Multidimensional Graph Pruning relies on LLM-as-a-Judge for overall functionality verification. To assess its reliability, we randomly sampled 50 accepted and 50 rejected samples for manual inspection. The results showed that all 50 accepted samples were indeed valid, while 41 of the 50 rejected samples were genuinely flawed. This suggests the LLM judge is highly reliable, ensuring that Multidimensional Graph Pruning effectively removes low-quality sample.
8.2. Computational Overhead Discussion